PLEASE HELP! I am not good with story problems and i am stump with this story equation. Here it is...
Steve, Theresa, and Christine play a round of golf together. Their combined score is 223. Steves score was 3 more than Christine's, and Theresa's score was 7 more than Steve's. What was each persons score?
Steve's score...
Theresa's score...
Christine's score...
Please help me on how to write the story problem to an equation.
THANK YOU for your help!

Answer by stanbon(75887)

Steve, Theresa, and Christine play a round of golf together. Their combined score is 223. Steves score was 3 more than Christine's, and Theresa's score was 7 more than Steve's. What was each persons score?
-------------------
Christine had the lowest score so let her score be "x".
Steve's score is "x+3"
Theresa's score = x+3+7 = "x+10"
----------------------
EQUATION:
x + (x+3) + (x+10) = 223
3x + 13 = 223
3x = 210
x = 70 (Christine's score)
x+3 = 73 (Steve's score)
x+10 = 80 (Theresa's score)
---------------------------------
Cheers,
Stan H.